OpenStudy (anonymous):

Help Finding the first four nonzero terms through power series expansion. y''-(cosx)y'-y=0 ; y(pi/2)=1 y'(pi/2)=1

OpenStudy (anonymous):

\[y''-(cosx)y'-y=0\]

OpenStudy (anonymous):

\[y(\frac{\pi}{2})=1,y'(\frac{\pi}{2})=1\]

OpenStudy (experimentx):

which crazy guy give this problem to you?

OpenStudy (anonymous):

my crazy differential equations prof

OpenStudy (experimentx):

good luck ... this is going to be terribly awful. first look into Fuchs-Frobenius theorem and hunt for singularities.

OpenStudy (anonymous):

Okay I have the solution, but am having a hard time writing it out http://www.math.ualberta.ca/~runde/files/sol201-8.pdf Scroll down to number 6

OpenStudy (experimentx):

then use power series for y(x) ...and expland cos(x) and then compare the coefficients of 'x'

OpenStudy (experimentx):

only 5 pages ??

OpenStudy (anonymous):

No it's several different problems, the question is number 6, it's worked out already

OpenStudy (anonymous):

An explanation would be beautiful.

OpenStudy (experimentx):

okay okay ... let's see what can I do.

OpenStudy (experimentx):

http://en.wikipedia.org/wiki/Fuchs's_theorem

OpenStudy (anonymous):

http://www.math.ualberta.ca/~runde/files/sol201-8.pdf pleaes look at number 6

OpenStudy (experimentx):

yes yes ... I am looking

OpenStudy (anonymous):

it's 5000 degrees in my country and the heat is getting to me im sorry.

OpenStudy (experimentx):

lol ... it's pretty hot here, but 00:19 in the morning!! so fine :) ... u go to alberta?? i thought it was in the north??

OpenStudy (anonymous):

I'm exaggerating, but please do explain if possible.

OpenStudy (experimentx):

haha i know ... by Fuchs theorem , at pi/2 p(x) is analytic.

OpenStudy (experimentx):

so it has a power series expansion of the form http://upload.wikimedia.org/math/6/b/3/6b34399a973da121c349586088ad1760.png

OpenStudy (experimentx):

where a is pi/2

OpenStudy (anonymous):

yes

OpenStudy (experimentx):

that 's' thing is useless there ... it's only useful when you have singularities in p(x) or q(x). so set s=0

OpenStudy (anonymous):

keep going

OpenStudy (experimentx):

 differentiate this thing once and twice ....